# SAM-CP: Marrying SAM with Composable Prompts for Versatile Segmentation

### v1

@@ Abstract
The Segment Anything model (SAM) has shown a generalized ability to group image pixels into patches, but applying it to semantic-aware segmentation still faces major challenges. This paper presents SAM-CP [^1] This work was done when Pengfei Chen ¡chenpengfei20@mails.ucas.ac.cn¿ was an intern at Huawei Inc.
* Corresponding author: Zhenjun Han ¡hanzhj@ucas.ac.cn¿.
, a simple approach that establishes two types of composable prompts beyond SAM and composes them for versatile segmentation. Specifically, given a set of classes (in texts) and a set of SAM patches, the Type-I prompt judges whether a SAM patch aligns with a text label, and the Type-II prompt judges whether two SAM patches with the same text label also belong to the same instance. To decrease the complexity in dealing with a large number of semantic classes and patches, we establish a unified framework that calculates the affinity between (semantic and instance) queries and SAM patches and merges patches with high affinity to the query. Experiments show that SAM-CP achieves semantic, instance, and panoptic segmentation in both open and closed domains. In particular, it achieves state-of-the-art performance in open-vocabulary segmentation. Our research offers a novel and generalized methodology for equipping vision foundation models like SAM with multi-grained semantic perception abilities.
@@ 1 Introduction
The past decade has witnessed a rapid development of vision-aware foundation models [40, 18, 50, 49, 36]. These models apply to a series of visual recognition tasks and serve as a building block for multimodal (e.g., vision-language) understanding. Recently, a powerful foundation model named the Segment Anything model (SAM) [24] has attracted a lot of attention. Pre-trained on a large corpus of images, SAM shows an impressive ability to group image pixels into patches and generalizes across various vision domains (e.g., medical images [21, 38, 22, 42], camouflaged images [9, 45], thermal images [6], etc.) as well as different downstream scenarios, (e.g., image editing [58, 51], 3D recognition [43, 48, 4], object tracking [13, 54], etc.).
Despite its success, there still exist major challenges in applying SAM to semantic-aware segmentation tasks including semantic, instance, or panoptic segmentation. We notice two lines of research in this direction. The first line (e.g., Grounded-SAM [41]) heavily relied on a standalone model (e.g., DINO [59] or Grounding-DINO [34]) to generate proposals and SAM was only used for refinement. This weakens the function of SAM as a foundation model. The second line (e.g., SSAM [5], Semantic-SAM [26], SAM-CLIP [47]) tried to assign a semantic label to each patch produced by SAM. However, in many scenarios, SAM may over-segment an instance into sub-patches, making it difficult to determine which patches belong to the same instance.
Figure 1: An illustration of how SAM-CP works at the idea level. Given an image and the patches produced by SAM, we first execute Prompt I to find the patches corresponding to any text label (in either closed or open domains), and then, if necessary, execute Prompt II to group the patches within each class into instances. In the upper part, the height of each bar corresponds to the probability that a patch belongs to a text label; in the lower part, two patches are connected by a solid line if they belong to the same instance. This figure is best viewed in color.
@@
We train SAM-CP on COCO [32] and ADE20K [61] and evaluate it on these two datasets as well as Cityscapes [14] for both open-vocabulary and closed-domain segmentation. Since SAM-CP is trained to understand text labels, it can easily adapt to unseen classes with a CLIP [40] text encoder. Extensive experiments demonstrate SAM-CP’s ability to cover semantic, instance, and panoptic segmentation with a single model. In particular, it reports state-of-the-art accuracy in open-vocabulary segmentation. Qualitative studies demonstrate that SAM-CP improves the semantic discriminativity of SAM’s features. Our research offers a new methodology to equip vision foundation models (e.g., SAM) with a solid and flexible ability for semantic recognition. We expect the proposed approach to gain stronger and more versatile abilities (e.g., for part segmentation) in the future, with the vision foundation models being upgraded and becoming more robust.
@@ 2 Related Work
In recent years, both the CV and MLP communities have witnessed a rapid development of foundation models. In particular, the vision foundation models have largely evolved from being simply pre-trained for image classification [16, 36, 49] to incorporating multimodal information [40, 17, 28, 2] and/or being pre-trained to deal with different tasks [23, 57, 50]. They improve the accuracy of various downstream visual recognition tasks including detection, segmentation, etc.
Recently, SAM [24] appeared as a foundation model for versatile segmentation. Pre-trained on a large corpus with billions of instances, SAM can segment an image into a set of basic patches. An important advantage of SAM lies in its ability to recognize images in different domains, yet a disadvantage lies in the lack of semantic labels on each patch. The community has been trying to adapt SAM to various scenarios, including adapting it on other image data (e.g., medical images [21, 38, 22, 42], camouflaged images [9, 45], remote sensing images [8, 60, 46], etc.), raising it to segment 3D objects [43, 48, 4], and using it as the pre-processing step of image editing [58, 51]. Among these efforts, one of the most challenging topics is to assign the SAM patches with semantic labels. Existing works involved calling other foundation models (e.g., CLIP [40]) for image tagging [5, 26, 47], applying SAM as a refinement stage after other detection and/or segmentation models [41], and generating other variants. However, in many scenarios, SAM may over-segment basic semantic units into sub-patches, which increases our burden of segmentation for specific purposes.
This paper focuses on establishing basic prompts beyond the segmentation results of SAM for versatile segmentation. This is related to a series of query-based algorithms for visual recognition, such as DETR [3] and the subsequent variants [62, 25, 33, 59, 34]. Meanwhile, we compute the affinity to determine the relationship between semantic units (e.g., queries and objects), which is related to a few prior work [1, 55, 35] that tried to compute the affinity between pixels and objects to perform segmentation. The idea was also inspired by ViRReq [44], a recent work that proposed decomposing complex visual recognition tasks into elementary units to ease annotation and optimization. Further, motivated by the previous open-vocabulary panoptic segmentation methods [56, 52, 10], we leverage a CLIP-based classifier to equip the model with an open-domain recognition ability.
@@ 3.1 Overview: Composite Prompts for Segmentation
The overall design of our approach is illustrated in Figure 1. The core idea is to establish two types of prompts beyond SAM [24], a recent vision foundation model that extracts patches from an input image as potential instances. By composing the output of different prompts, the SAM patches are labelled and/or combined into semantic regions and/or instances, and thus versatile segmentation tasks can be performed. We name the approach SAM-CP, where ‘CP’ stands for composable prompts.
Mathematically, let the input image be $\mathbf{I}$. SAM extracts a number of patches, $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{N}\}$, where $N$ is the number of patches and $\mathbf{P}_{n}$ is the $n$-th patch which is represented as a binary mask of the same shape as the input image. Although SAM is robust across various vision domains, it does not offer a semantic label to each patch and, sometimes, an instance (e.g., a person) may be over-segmented into multiple patches. We design the following two types of prompts.
@@ 3.2.1 Patch Encoder
We first extract visual features from the patches. For each patch $\mathbf{P}_{n}$, we apply a regular backbone (e.g., ResNet50 [20] or Swin-L [36]) equipped with an RoIAlign [19] operator to obtain a basic feature vector $\tilde{\mathbf{f}}_{n}$. We also design a MaskRoI operator to extract more accurate visual features by masking out the background areas. All these features are propagated through a multi-layer perceptron (MLP) and fed into $\omega$ multi-head self-attention layers, where $\omega$ is set to 6 throughout this paper. The feature vector corresponding to $\mathbf{P}_{n}$ is denoted as $\mathbf{f}_{n}$.
@@ 3.2.2 Unified Affinity Decoder
Queries. The queries are of a similar form as in DETR [3]. Differently, we establish two types of queries for semantic and instance segmentation, respectively. (1) For each text label $\mathbf{T}_{c}$ where $c\in\{1,2,\ldots,C\}$ is the class index, we use the language branch of a vision-language model (e.g., CLIP [40]) to convert it into a query vector, $\mathbf{e}_{c}^{\mathrm{S}}$, where the superscript ‘$\mathrm{S}$’ stands for ‘semantic’. (2) We also create $N$ instance queries (i.e., assuming that each patch may correspond to an instance, abbr. patch-as-query, PasQ) and initialize them using the visual features and position embeddings of the patches, i.e., $\mathbf{e}_{n}^{\mathrm{I}}\equiv\mathbf{f}_{n}$, where the superscript ‘$\mathrm{I}$’ stands for ‘instance’. In what follows, we denote both types of queries as $\mathbf{Q}_{m}$, $m=1,2,\ldots,M$ where $M$ is the number of queries.
@@ 3.2.3 Label Assignment and Supervision
Semantic-level supervision. We first build a vision-language classifier upon the semantic queries. Following GLIP [29], the score $S_{m,c}^{\mathrm{cls}}$ of the $m$-th query at the $c$-th class is determined by $\hat{\mathbf{Q}}_{m}$ and $\hat{\mathbf{e}}_{c}$, the quantities produced by linearly normalizing $\mathbf{Q}_{m}$ and $\mathbf{e}_{c}$ into $[0,1]$. The classification loss [31] $\mathcal{L}_{\mathrm{cls}}$ is computed upon $S_{m,c}$ in Equation (1).
@@
In $S_{m,c}^{\mathrm{cls}}$, $s$ is a learnable scaling factor, and $b$ is a bias parameter. In $\mathcal{L}_{\mathrm{cls}}$, $\mathrm{FL}(\cdot,\cdot)$ is the focal loss, $\sigma(\cdot)$ is the sigmoid activation function, $c_{m}^{\star}$ is the ground-truth class label for the $m$-th query (we explain how to compute $c_{m}^{\star}$ in the supplementary material), and $\mathbb{I}_{[\cdot]}$ the indicator function which takes $1$ if the statement is true and 00 otherwise.
Instance-level supervision. At the end of affinity propagation, each instance query corresponds to a binary segmentation mask. Let the ground truth contain $K$ instances. We first establish a matching matrix $\mathbf{G}$ (sized $K\times N$) by computing the box-level IoP (intersection-over-patch) and mask-level IoP values between each pair of predicted and true instances; they are considered matched if both IoP values are greater than a pre-defined hyper-parameter $\tau$, i.e., $G_{k,n}=\mathbb{I}_{[\min(\mathrm{IoP}_{\mathrm{box}},\mathrm{IoP}_{\mathrm{mask}})>\tau]}$, where $\tau=0.8$ throughout this paper. If no patches are assigned to an object, the patches with an IoU value of at least $0.5$ will be chosen as candidates of low-quality matching. Based on the matching matrix $\mathbf{G}$, we compute the ground-truth affinity matrix $\mathbf{B}$ (sized $M\times N$, same as $\mathbf{A}$). For each $m$, we first determine if any ground-truth instance (indexed $k_{m}$) matches the $m$-th query (see the next part for details). If yes, $\mathbf{B}_{m}=\mathbf{G}_{k_{m}}$ (i.e., the $k_{m}$-th row of $\mathbf{G}$ is copied to the $m$-th row of $\mathbf{B}$); otherwise, $\mathbf{B}_{m}\equiv\mathbf{0}$. Then, following [11, 27], we compute the mask focal loss $\mathcal{L}_{\mathrm{mfl}}$ and the Dice loss $\mathcal{L}_{\mathrm{dice}}$:
@@
  • For an instance (Type-II) queries, we follow the DETR series [3, 62, 59] to apply the Hungarian algorithm to find the best matches between the queries and the ground-truth instances. Differently, we compute the matching cost using more metrics, including the classification loss (cls), the mask focal loss (mfl), the Dice loss (dice), the bounding-box cost (bbox), and the gIoU cost (giou). We will show in experiments that all these components contribute to better segmentation results. After the matching is done, we obtain the index $k_{m}$ for the $m$-th query ($k_{m}$ can be null, in which situation the query is ignored), and assign $\mathbf{B}_{m}=\mathbf{G}_{k_{m}}$. We set $\varepsilon_{m}$ to $1$ for the positive embeddings and 00 otherwise.
The overall loss function. The overall loss is defined as $\mathcal{L}_{\mathrm{all}}=\lambda_{\mathrm{cls}}\cdot\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{mfl}}\cdot\mathcal{L}_{\mathrm{mfl}}+\lambda_{\mathrm{dice}}\cdot\mathcal{L}_{\mathrm{dice}}$, where the loss coefficients are $\lambda_{\mathrm{cls}}=2$, $\lambda_{\mathrm{mfl}}=1$, and $\lambda_{\mathrm{dice}}=1$. We apply the denoising strategy in DINO [59] to improve the training performance. The details are provided in Appendix A.3.
3.3 Inference
The inference procedure varies slightly between closed-set and open-vocabulary segmentation. In a closed domain, the logit $\mathbf{S}^{\mathrm{cls}}$ in Equation (1) is used for classification. The normalized and quantized affinity matrix $\mathbf{A}$ is used for patch merging. For semantic segmentation, we refer to the rows corresponding to $C$ classes and merge all patches surpassing a pre-defined threshold. For instance segmentation, we look up the non-empty instance rows of $\mathbf{A}$, each of which corresponds to an instance. Panoptic segmentation is achieved by combining semantic and instance segmentation results. In an open domain, we complement the logit matrix $\mathbf{S}^{\mathrm{cls}}$ (sized $M\times C$) with a CLIP-based logit matrix, $\mathbf{S}^{\mathrm{CLIP}}$. The remaining part is the same as in a closed domain. To calculate $\mathrm{S}^{\mathrm{CLIP}}$, we follow FC-CLIP [56] to extract CLIP features using mask pooling on the predicted masks. Then, the feature $\mathbf{S}^{\mathrm{CLIP}}$ is obtained by calculating the similarity between the CLIP feature and $\hat{\mathbf{e}}_{c}$. The final class score is computed as $\mathbf{S}^{\mathrm{ov}}=(\sigma(\mathbf{S}^{\mathrm{cls}}))^{(1-\kappa)}+(\mathrm{softmax}(\mathbf{S}^{\mathrm{CLIP}}))^{\kappa}$, where $\kappa=0.4$ is a coefficient balancing the closed-domain and open-vocabulary class scores.

### v2

@@ Abstract
The Segment Anything model (SAM) has shown a generalized ability to group image pixels into patches, but applying it to semantic-aware segmentation still faces major challenges. This paper presents SAM-CP, a simple approach that establishes two types of composable prompts beyond SAM and composes them for versatile segmentation. Specifically, given a set of classes (in texts) and a set of SAM patches, the Type-I prompt judges whether a SAM patch aligns with a text label, and the Type-II prompt judges whether two SAM patches with the same text label also belong to the same instance. To decrease the complexity in dealing with a large number of semantic classes and patches, we establish a unified framework that calculates the affinity between (semantic and instance) queries and SAM patches, and then merges patches with high affinity to the query. Experiments show that SAM-CP achieves semantic, instance, and panoptic segmentation in both open and closed domains. In particular, it achieves the state-of-the-art performance in open-vocabulary segmentation. Our research offers a novel and generalized methodology for equipping vision foundation models like SAM with multi-grained semantic perception abilities. Codes are released on github.com/ucas-vg/SAM-CP.
@@ 1 Introduction
The past decade has witnessed a rapid development of vision-aware foundation models Radford et al. (2021); He et al. (2022); Wang et al. (2023c; 2021); Liu et al. (2021). These models apply to a series of visual recognition tasks and serve as a building block for multimodal (e.g., vision-language) understanding. Recently, a powerful foundation model named the Segment Anything model (SAM) Kirillov et al. (2023) has attracted a lot of attention. Pre-trained on a large corpus of images, SAM shows an impressive ability to group image pixels into patches and generalizes across various vision domains (e.g., medical images He et al. (2023); Ma et al. (2024); Hu & Li (2023); Roy et al. (2023), camouflaged images Chen et al. (2023b); Tang et al. (2023b), thermal images Chen & Bai (2023), etc.) as well as different downstream scenarios, (e.g., image editing Yu et al. (2023c); Xie et al. (2023), 3D recognition Shvets et al. (2024); Wang et al. (2023b); Cen et al. (2023), object tracking Cheng et al. (2023); Yang et al. (2023), etc.).
Despite its success, there still exist major challenges in applying SAM to semantic-aware segmentation tasks including semantic, instance, or panoptic segmentation. We notice two lines of research in this direction. The first line (e.g., Grounded-SAM Ren et al. (2024)) heavily relied on a standalone model (e.g., DINO Zhang et al. (2023a) or Grounding-DINO Liu et al. (2024)) to generate proposals and SAM was only used for refinement. This weakens the function of SAM as a foundation model. The second line (e.g., SSAM Chen et al. (2023a), Semantic-SAM Li et al. (2023a), SAM-CLIP Wang & Vasu (2023)) tried to assign a semantic label to each patch produced by SAM. However, in many scenarios, SAM may over-segment an instance into sub-patches, making it difficult to determine which patches belong to the same instance.
Figure 1: An illustration of how SAM-CP works at the idea level. Given an image and the patches produced by SAM, we first execute Prompt I to find the patches corresponding to any text label (in either closed or open domains), and then, if necessary, execute Prompt II to group the patches within each class into instances. In the upper part, the height of each bar corresponds to the probability that a patch belongs to a text label (yellow, green, blue and red for ‘sand’, ‘kite’, ‘sky’ and ‘person’); in the lower part, two patches are connected by a solid line if they belong to the same instance (purple for ‘person-1’ and orange for ‘person-2’). This figure is best viewed in color.
@@
We train SAM-CP on COCO Lin & Maire (2014) and ADE20K Zhou et al. (2017) and evaluate it on these two datasets as well as Cityscapes Cordts et al. (2016) for both open-vocabulary and closed-domain segmentation. Since SAM-CP is trained to understand text labels, it can easily adapt to unseen classes with a CLIP Radford et al. (2021) text encoder. Extensive experiments demonstrate SAM-CP’s ability to cover semantic, instance, and panoptic segmentation with a single model. In particular, it reports state-of-the-art accuracy in open-vocabulary segmentation. Qualitative studies demonstrate that SAM-CP improves the semantic discriminativity of SAM’s features. Our research offers a new methodology to equip vision foundation models (e.g., SAM) with a solid and flexible ability for semantic recognition. We expect the proposed approach to gain stronger and more versatile abilities in the future, with the vision foundation models being upgraded and becoming more robust.
@@ 2 Related Work
In recent years, both the CV and NLP communities have witnessed a rapid development of foundation models. In particular, the vision foundation models have largely evolved from being simply pre-trained for image classification Dosovitskiy et al. (2021); Liu et al. (2021); Wang et al. (2021) to incorporating multimodal information Radford et al. (2021); Fang et al. (2023); Li et al. (2023c); Alayrac et al. (2022) and/or being pre-trained to deal with different tasks Huang et al. (2023); Yu et al. (2023b); Wang et al. (2023c). They improve the accuracy of various downstream visual recognition tasks including detection, segmentation, etc.
Recently, SAM Kirillov et al. (2023) appeared as a foundation model for versatile segmentation. Pre-trained on a large corpus with billions of instances, SAM can segment an image into a set of basic patches without tuning. An important advantage of SAM lies in its ability to recognize images in different domains, yet a disadvantage lies in the lack of semantic labels on each patch. The community has been trying to adapt SAM to various scenarios, including adapting it on other image data (e.g., medical images He et al. (2023); Ma et al. (2024); Hu & Li (2023); Roy et al. (2023), camouflaged images Chen et al. (2023b); Tang et al. (2023b), remote sensing images Chen et al. (2024a); Zhang et al. (2023c); Wang et al. (2023a), etc.), raising it to segment 3D objects Shvets et al. (2024); Wang et al. (2023b); Cen et al. (2023), and using it as the pre-processing step of image editing Yu et al. (2023c); Xie et al. (2023). Among these efforts, one of the most challenging topics is to assign the SAM patches with semantic labels. Existing works involved calling other foundation models (e.g., CLIP Radford et al. (2021)) for image tagging Chen et al. (2023a); Li et al. (2023a); Wang & Vasu (2023); Wei et al. (2024), applying SAM as a refinement stage after other detection and/or segmentation models Ren et al. (2024), and generating other variants. However, in many scenarios, SAM may over-segment basic semantic units into sub-patches, which increases our burden of segmentation for specific purposes.
This paper focuses on establishing basic prompts beyond the segmentation results of SAM for versatile segmentation. This is related to a series of query-based algorithms for visual recognition, such as DETR Carion et al. (2020) and the subsequent variants Zhu et al. (2021); Li et al. (2022a); Liu et al. (2022); Zhang et al. (2023a); Liu et al. (2024). Meanwhile, we compute the affinity to determine the relationship between semantic units (e.g., queries and objects), which is related to a few prior work Ahn & Kwak (2018); Yu et al. (2020); Liu et al. (2017) that tried to compute the affinity between pixels and objects to perform segmentation. The idea was also inspired by ViRReq Tang et al. (2023a), a recent work that proposed decomposing complex visual recognition tasks into elementary units to ease annotation and optimization. Further, motivated by the previous open-domain panoptic segmentation methods Yu et al. (2023a); Xu et al. (2023a); Chen et al. (2023c), we leverage a CLIP-based classifier to equip the model with an open-domain recognition ability.
@@ 3.1 Overview: Composite Prompts for Segmentation
The overall design of our approach is illustrated in Figure 1. The core idea is to establish two types of prompts beyond SAM Kirillov et al. (2023), a recent vision foundation model that extracts patches from an input image as potential instances. By composing the output of different prompts, the SAM patches are labelled and/or combined into semantic regions and/or instances, and thus versatile segmentation tasks can be performed. We name the approach SAM-CP, where ‘CP’ stands for composable prompts.
Mathematically, let the input image be $\mathbf{I}$. SAM extracts a number of patches, $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{N}\}$, under segment-everything mode. $N$ is the number of patches and $\mathbf{P}_{n}$ is the $n$-th patch which is represented as a binary mask of the same shape as the input image. Although SAM is robust across various vision domains, it does not offer a category to each patch and, sometimes, an instance (e.g., a person) may be over-segmented into multiple patches. We design the following two types of prompts.
@@ 3.2.1 Patch Encoder
We first extract visual features from the patches. For each patch $\mathbf{P}_{n}$, we apply a regular backbone (e.g., ResNet50 He et al. (2016) or Swin-L Liu et al. (2021)) equipped with an RoIAlign He et al. (2017) operator to obtain a basic feature vector $\tilde{\mathbf{f}}_{n}$. We also design a MaskRoI operator to extract more accurate visual features by masking out the background areas. All these features are propagated through a multi-layer perceptron (MLP) and fed into $\omega$ multi-head self-attention layers, where $\omega$ is set to 6 throughout this paper. The feature vector corresponding to $\mathbf{P}_{n}$ is denoted as $\mathbf{f}_{n}$.
@@ 3.2.2 Unified Affinity Decoder
Queries. The queries are of a similar form as in DETR Carion et al. (2020). Differently, we establish two types of queries for semantic and instance segmentation, respectively. (1) For each text label $\mathbf{T}_{c}$ where $c\in\{1,2,\ldots,C\}$ is the class index, we use the language branch of a vision-language model (e.g., CLIP Radford et al. (2021)) to convert it into a query vector, $\mathbf{e}_{c}^{\mathrm{S}}$, where the superscript ‘$\mathrm{S}$’ stands for ‘semantic’. (2) We also create $N$ instance queries (i.e., assuming that each patch may correspond to an instance, abbr. patch-as-query, PasQ) and initialize them using the visual features and position embeddings of the patches, i.e., $\mathbf{e}_{n}^{\mathrm{I}}\equiv\mathbf{f}_{n}$, where the superscript ‘$\mathrm{I}$’ stands for ‘instance’. We denote both types of queries as $\mathbf{Q}_{m}$, $m=1,2,\ldots,M$ where $M$ is the number of queries.
@@ 3.2.3 Label Assignment and Supervision
Semantic-level supervision. We first build a vision-language classifier upon the semantic queries. Following GLIP Li et al. (2022b), the score $S_{m,c}^{\mathrm{cls}}$ of the $m$-th query at the $c$-th class is determined by $\hat{\mathbf{Q}}_{m}$ and $\hat{\mathbf{e}}_{c}$, the quantities produced by linearly normalizing $\mathbf{Q}_{m}$ and $\mathbf{e}_{c}$ into $[0,1]$. The classification loss Lin et al. (2017b) $\mathcal{L}_{\mathrm{cls}}$ is computed upon $S_{m,c}$ in Equation 1.
@@
In $S_{m,c}^{\mathrm{cls}}$, $s$ is a learnable scaling factor, and $b$ is a bias parameter. In $\mathcal{L}_{\mathrm{cls}}$, $\mathrm{FL}(\cdot,\cdot)$ is the focal loss, $\sigma(\cdot)$ is the sigmoid activation function, $c_{m}^{\star}$ is the ground-truth class label for the $m$-th query (we explain how to compute $c_{m}^{\star}$ in the Appendix A.2), and $\mathbb{I}_{[\cdot]}$ the indicator function which takes $1$ if the statement is true and 00 otherwise. Note that, if we only perform Type-I prompts, the category text is not necessary, and we only need to perform a binary prediction for each text query to judge whether the category exists in the image. This mechanism was designed for Type-II prompts. An important technical contribution of our work is to unify Type-I and Type-II prompts into one framework so that the training and inference costs are reduced.
Instance-level supervision. At the end of affinity propagation, each instance query corresponds to a binary segmentation mask. Let the ground truth contain $K$ instances. We first establish a matching matrix $\mathbf{G}$ (sized $K\times N$) by computing the box-level IoP (intersection-over-patch) and mask-level IoP values between each pair of predicted and true instances; they are considered matched if both IoP values are greater than a pre-defined hyper-parameter $\tau$, i.e., $G_{k,n}=\mathbb{I}_{[\min(\mathrm{IoP}_{\mathrm{box}},\mathrm{IoP}_{\mathrm{mask}})>\tau]}$, where $\tau=0.8$ throughout this paper. If no patches are assigned to an object, the patches with an IoU value of at least $0.5$ will be chosen as candidates of low-quality matching. Based on the matching matrix $\mathbf{G}$, we compute the ground-truth affinity matrix $\mathbf{B}$ (sized $M\times N$, same as $\mathbf{A}$). For each $m$, we first determine if any ground-truth instance (indexed $k_{m}$) matches the $m$-th query (see the next part for details). If yes, $\mathbf{B}_{m}=\mathbf{G}_{k_{m}}$ (i.e., the $k_{m}$-th row of $\mathbf{G}$ is copied to the $m$-th row of $\mathbf{B}$); otherwise, $\mathbf{B}_{m}\equiv\mathbf{0}$. Then, following Cheng et al. (2022); Li et al. (2023b), we compute the mask focal loss $\mathcal{L}_{\mathrm{mfl}}$ and the Dice loss $\mathcal{L}_{\mathrm{dice}}$:
@@
  • For an instance (Type-II) queries, we follow the DETR series Carion et al. (2020); Zhu et al. (2021); Zhang et al. (2023a) to apply the Hungarian algorithm to find the best matches between the queries and the ground-truth instances. Differently, we compute the matching cost using more metrics, including the classification loss (cls), the mask focal loss (mfl), the Dice loss (dice), the bounding-box cost (bbox), and the gIoU cost (giou). We will show in experiments that all these components contribute to better segmentation results. After the matching is done, we obtain the index $k_{m}$ for the $m$-th query ($k_{m}$ can be null, in which situation the query is ignored), and assign $\mathbf{B}_{m}=\mathbf{G}_{k_{m}}$. We set $\varepsilon_{m}$ to $1$ for the positive embeddings and 00 otherwise.
The overall loss function. The overall loss is defined as $\mathcal{L}_{\mathrm{all}}=\lambda_{\mathrm{cls}}\cdot\mathcal{L}_{\mathrm{cls}}+\lambda_{\mathrm{mfl}}\cdot\mathcal{L}_{\mathrm{mfl}}+\lambda_{\mathrm{dice}}\cdot\mathcal{L}_{\mathrm{dice}}$, where the loss coefficients are $\lambda_{\mathrm{cls}}=2$, $\lambda_{\mathrm{mfl}}=1$, and $\lambda_{\mathrm{dice}}=1$. We apply the denoising strategy in DINO Zhang et al. (2023a) to improve the training performance.